Circular No. 7981

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002ez, 2002fv, 2002fw, 2002fx, 2002fy, 2002fz, 2002ga

M. Giavalisco and A. Riess, Space Telescope Science Institute, on behalf of the data searchers of the ‘Great Observatories Origins Deep Survey (GOODS) Treasury Team’ (M. Giavalisco, S. Casertano, T. Dahlen, M. Dickinson, H. Ferguson, R. Hook, R. Idzi, A. Koekemoer, B. Mobasher, L. A. Moustakas, S. Ravindranath) and the Hubble Higher-z Supernova Team (A. Riess, L. Strolger, J. Tonry, P. Challis), report the discovery of six apparent new supernovae on CCD images taken in the interval Sept. 19–22 with the Advanced Camera for Surveys (ACS) on the Hubble Space Telescope during the GOODS observations of the Chandra Deep Field South. The apparent supernovae were detected by subtracting two groups of ACS images of the same fields, the first taken during July 31–Aug. 4 and the second durng Sept. 19–22. Each new object was detected independently in each of four separate z-band (F850lp) exposures and two i-band (F775W) exposures taken during Sept. 19–22, with five of the six objects also detected in two v-band (F606W) exposures. The morphology of each transient is consistent with that of the ACS Point Spread Function (FWHM = 0′′.1). The lack of detectable motion (< 0′′.05) relative to other sources in the image over a time span of ∼ 45 min eliminates the possibility that these are small bodies in the solar system. The following table summarizes the findings. The z-band magnitudes below are relative to Vega, and uncertainties are typically 0.05–0.1 mag (random), with a systematic zero-point uncertainty of 0.1 mag. SNe 2002fv and 2002fy have no apparent host galaxies.

SN  2002 UT  α_{2000}  δ_{2000}  z  Offset
2002fv Sept. 19.58 3^{h}32^{m}19.^{s}22 −27^{o}49′34′′.0 25.2  —
2002fw  19.86 3 32 37.52 −27 46 46.6 24.6 0′′.51 W, 0′′.21 N
2002fx  20.84 3 32 06.80 −27 44 34.4 25.5 0′′.09 W, 0′′.04 S
2002fy 20.99 3 32 18.12 −27 41 55.6 24.9 —
2002fz  21.60 3 32 48.54 −27 54 17.6 24.9 1′′.30 E, 1′′.66 S
2002ga  22.52 3 32 32.62 −27 53 16.7 24.9 0′′.21 E, 0′′.08 S

Additional z-band magnitude estimates: SN 2002fv, Aug. 4.19 UT, [28.0; SN 2002fw, Aug. 1.39, [28.0; SN 2002fx, Aug. 4.33, [28.0; SN 2002fy, Aug. 3.92, 23.8; SN 2002fz, Aug. 1.12, 25.5; SN 2002ga, Aug. 2.26, 26.3.

Corrigendum. On IAUC 7967, the offset of SN 2002ez from the center of its apparent host galaxy should read 4′′ E, 24′′ S.

2002 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green